### v1

@@ 5.1.1 Phase 1 (Registering neighbour agents, adjacent edges and port numbers)
With each node of $G$ hosting a distinct agent, each agent first visits and registers its list of neighbours. Due to a lack of movement synchronisation, the agents follow the exact protocol as described in Phase 1 of Section 4 to register the neighbours, the ports which lead to the respective neighbours and the respective edges. To avoid inconsistency, the agents maintain the record of the edges in a specific order. An agent $r_{i}$ only stores the record of an edge $(r_{i},r_{j})$ only if $r_{j}$ is a neighbour of $r_{i}$ and $r_{j}.ID>r_{i}.ID$. This gives a total ordering on the edges. This phase takes $O(\Delta\cdot\log\lambda)$ rounds to complete. Now the agents begin the next phase of calculating $support$ for each edge. With the execution of Phase 1, $edge\_set$ gets recorded.
@@ 5.1.2 Phase 2 ($support$ calculation and initial $h$ initialization)
The agents now calculate the support for each edge. To do this, the agents use a similar methodology to the one used in Phase 2 of Section 4. The agents go to each of their neighbours once again. Let us assume, at some point of time during the algorithm, an agent $r_{i}$ meets its neighbour $r_{j}$. $r_{i}$ then checks the number of common neighbours with $r_{j}$ and records the count into $r_{i}.support(r_{i},r_{j})$ (we have assumed that $r_{j}.ID>r_{i}.ID$; otherwise if $r_{j}.ID<r_{i}.ID$, then the edge $(r_{i},r_{j})$ (simultaneously, $r_{i}.support(r_{i},r_{j})$) does not register in $r_{i}$’s records and in such case, the agent $r_{i}$ does nothing. However, in such case, the edge does register in $r_{j}$’s record, and $r_{j}$ then does the needful in that respect.). Support calculation takes $O(\Delta\log\lambda)$ rounds, after which, each agent receives its $support$ number. With the execution of this, $h$ value for each edge in $r_{i}.edge\_set$ gets initialized to the support of the particular edge.
@@ 5.1.3 Phase 3 (Iterative $h-index$ up-gradation)
With each edge receiving its initial $support$ value, the algorithm now updates $h$ for each edge through multiple iterations of this phase. The phase iterates till the $change$ value in any of the agents $r_{i}$ remains 00 and terminates once the value of $change$ in each and every agent becomes $1$. In each running phase, each agent $r_{i}$ does the following in parallel.
@@
    Checks if any edge is scheduled for further up-gradation in its $edge\_set$. If no, set $r_{i}.change\leftarrow 1$ and stops executing the algorithm but remains active. $r_{i}$ provides any information when required to other agents.
@@
    For each edge $e$ in $edge\_set$ which is scheduled for up-gradation, $L$ stores $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ in $G$. To obtain $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must communicate with its neighbouring agent. For example, consider the edge $(r_{i},r_{j})$ with $r_{i}<r_{j}$, stored in $r_{i}.edge\_set$. Also, consider another agent $r_{p}(r_{p}<r_{i}<r_{j})$ stationed on another node such that $(r_{i},r_{j},r_{p})$ forms a triangle. Let $e=(r_{i},r_{j}),e^{\prime}=(r_{p},r_{i}),e^{\prime\prime}=(r_{p},r_{j})$. Now for $r_{i}$ to calculate $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must visit $r_{p}$. Since the movement of the agents is not centralised, $r_{i}$ must execute "Know the Neighbourhood" (Phase 1, Sec. 4.1) which takes additional $O(\Delta\log\lambda)$ rounds, in order to guarantee that $r_{i}$ meets $r_{p}$. Along with the set $L$, the set $N$ is also updated. Let $e=(r_{i},r_{j})$. For each triangle $(e,e^{\prime},e^{\prime\prime})$ the agents add $e^{\prime},e^{\prime\prime}$ to their respective $N$ sets. Now, since each edge, $e^{\prime}$ or $e^{\prime\prime}$ is associated with a unique agent, an invocation of "Know the Neighbourhood" (Phase 1, Sec. 4.1) is required to inform the correct agent associated with a particular edge so that the agent can schedule the particular edge in the next step if required.
@@
    Checking for Termination: Each agent, after having evaluated the $change$ value for the current phase, communicates with each of the other agents to check, if the algorithm is ready for termination. The agent sends the $change$ values to every agent in multiple rounds. To communicate, each agent starts by visiting its immediate neighbours. In the first round, each agent sends the $change$ to its neighbour. In the second round, the same change value is propagated to the 2-hop neighbour and so on. By the end of $D$ rounds, each agent receives every changed value from every other agent. Now, the agents calculate the binary product of all the $change$ values received. If the binary product of the $change$ value results in $1$ (all the agents must get the same binary product value), the agents terminate the algorithm. Otherwise, the agents rerun Phase 3 again, till the product becomes $1$.
Lemma 7.
Algorithm 5.1 takes $O(m\Delta D\log\lambda)$ rounds to terminate
@@
Phase 1 and Phase 2 takes $O(\Delta\log\lambda)$ rounds, as shown in Lemma 2 and Lemma 3. In each execution of Phase 3, the agent needs to communicate with its neighbour to update its $trussness(h)$ values taking $O(\Delta\log\lambda)$ rounds. Finally, to check for termination, each agent propagates the $change$ value to every other agent, needing $O(D\Delta\log\lambda)$ rounds.
At worst, the algorithm may update the $h-index$ value of one edge during each execution of Phase 3. Also, the up-gradation of the $h-index$ value of an edge is equivalent to the fact that the triangle count ($support)$ of that particular edge (and possibly some other edges) in $G$ has been decreased. Therefore, each time an $h-index$ of an edge is updated (it can only decrease), the graph $G$ loses at least one edge. Therefore, the algorithm may need to execute Phase 3, $m$ times before termination. Therefore Algorithm 5.1 takes $O(\Delta\log\lambda)+m\cdot(O(\Delta\log\lambda)+O(D\Delta\log\lambda))=O(m\Delta D\log\lambda)$ rounds to execute. ∎
@@
Let $G$ be an $n$-node arbitrary, simple, connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest agent ID $\lambda\in[0,n^{c}]$, where $c$ is constant, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, the Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\cdot\log n)$ bits of memory per agent.
@@ 5.2 Triangle Centrality
where, where $N(v)$ is the neighborhood set of $v$, $N_{\mathbf{T}}(v)$ is the set of neighbors that are in triangles with $v$, and $N^{+}_{\mathbf{T}}(v)$ is the closed set that includes $v$. $\mathbf{T}(v)$ and $\mathbf{T}(G)$ denote the respective triangle counts based on $v$ and total triangle count in $G$. Mathematically, $N(v)=\{u:(u,v)\in E(G)\}$, $N_{\mathbf{T}}(v)=\{u\in N(v):N(u)\cap N(v)\neq\phi\}$ and $N^{+}_{\mathbf{T}}(v)=\{v\}\cup N_{\mathbf{T}}(v)$. Now, we outline our algorithm that computes $TC(v)$ for a node $v\in G$ using the mobile agents that runs in these $3$ phases:
@@
    Computing $\mathbf{T}(v)$ and $\mathbf{T}(G)$ We use the triangle counting algorithm for mobile agents described in Section 4 to evaluate the number of triangles involving the node $v$, $\mathbf{T}(v)$ and the total number of triangles in $G$, $\mathbf{T}(G)$. The execution of this phase takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda)=O(D\Delta\log\lambda)$ rounds [Lemma 3,6]. With the completion of this phase, the agents begin the next phase. Note that if the count of the total number of triangles is available apriori, only $T(v)$ needs to evaluated.
@@
    Computing $N(v)$, $N^{+}_{\mathbf{T}}(v)$ and $N_{\mathbf{T}}(v)$ $N(v)$, the neighbours of $v$ can be recorded in $O(\Delta\log\lambda)$ rounds by the agents (Lemma 2). After we allow the agents to record their neighbourhood in the first $O(\Delta\log\lambda)$ rounds, in the subsequent $O(\Delta\log\lambda)$ rounds, the agents can now meet their neighbours once again, with this time, learning about $N(u)$ for each neighbour agent $u$ of $v$. Now the agent from $v$ communicates with each $u\in N(v)$ to evaluate the set $N_{\mathbf{T}}(v)$. Similarly, the agents can also build the set $N^{+}_{\mathbf{T}}(v)$. Now, as the agent meets with each of its neighbours $u\in N(v)$, it cumulatively evaluates the sums $s=\sum_{u\in N^{+}_{\mathbf{T}}(v)}$ and $x=\sum_{u\in N(v)_{\mathbf{T}}(v)}$ simultaneously. Finally the sum $\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}=s-x+\mathbf{T}(v)$ is evaluated. This step takes $O(\Delta\log\lambda)$ rounds.

### v2

@@ 5.1.1 Phase 1 (Registering neighbour agents, adjacent edges and port numbers)
With each node of $G$ hosting a distinct agent, each agent first visits and registers its list of neighbours. Due to a lack of movement synchronisation, the agents follow the exact protocol as described in Phase 1 of Section 4 to register the neighbours, the ports that lead to the respective neighbours and the respective edges. To avoid inconsistency, the agents maintain the record of the edges in a specific order. An agent $r_{i}$ only stores the record of an edge $(r_{i},r_{j})$ only if $r_{j}$ is a neighbour of $r_{i}$ and $r_{j}.ID>r_{i}.ID$. This gives a total ordering on the edges. This phase takes $O(\Delta\cdot\log\lambda)$ rounds to complete. Now the agents begin the next phase of calculating $support$ for each edge. With the execution of Phase 1, $edge\_set$ gets recorded.
@@ 5.1.2 Phase 2 ($support$ calculation and initial $h$ initialization)
The agents now calculate the support for each edge. To do this, the agents use a similar methodology to the one used in Phase 2 of Section 4. The agents go to each of their neighbours once again. Let us assume, at some point of time during the algorithm, an agent $r_{i}$ meets its neighbour $r_{j}$. $r_{i}$ then checks the number of common neighbours with $r_{j}$ and records the count into $r_{i}.support(r_{i},r_{j})$ (we have assumed that $r_{j}.ID>r_{i}.ID$; otherwise if $r_{j}.ID<r_{i}.ID$, then the edge $(r_{i},r_{j})$ (simultaneously, $r_{i}.support(r_{i},r_{j})$) does not register in $r_{i}$’s records and in such a case, the agent $r_{i}$ does nothing. However, in such a case, the edge does register in $r_{j}$’s record, and $r_{j}$ then does the needful in that respect.). Support calculation takes $O(\Delta\log\lambda)$ rounds, after which, each agent receives its $support$ number. With the execution of this, the $h$ value for each edge in $r_{i}.edge\_set$ gets initialized to the support of the particular edge.
@@ 5.1.3 Phase 3 (Iterative $h-index$ up-gradation)
With each edge receiving its initial $support$ value, the algorithm now updates $h$ for each edge through multiple iterations of this phase. The phase iterates till the $change$ value in any of the agents $r_{i}$ remains 00 and terminates once the value of $change$ in each and every agent becomes $1$. In each phase, each agent $r_{i}$ does the following in parallel.
@@
    Checks if any edge is scheduled for further up-gradation in its $edge\_set$. If no, set $r_{i}.change\leftarrow 1$ and stop executing the algorithm but remain active. $r_{i}$ provides any information when required to other agents.
@@
    For each edge $e$ in $edge\_set$ which is scheduled for up-gradation, $L$ stores $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ in $G$. To obtain $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must communicate with its neighbouring agent. For example, consider the edge $(r_{i},r_{j})$ with $r_{i}<r_{j}$, stored in $r_{i}.edge\_set$. Also, consider another agent $r_{p}(r_{p}<r_{i}<r_{j})$ stationed on another node such that $(r_{i},r_{j},r_{p})$ forms a triangle. Let $e=(r_{i},r_{j}),e^{\prime}=(r_{p},r_{i}),e^{\prime\prime}=(r_{p},r_{j})$. Now for $r_{i}$ to calculate $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must visit $r_{p}$. Since the movement of the agents is not centralised, $r_{i}$ must execute ”Neighbourhood Discovery” (Phase 1, Sec. 4.1) which takes additional $O(\Delta\log\lambda)$ rounds, in order to guarantee that $r_{i}$ meets $r_{p}$. Along with the set $L$, the set $N$ is also updated. Let $e=(r_{i},r_{j})$. For each triangle $(e,e^{\prime},e^{\prime\prime})$ the agents add $e^{\prime},e^{\prime\prime}$ to their respective $N$ sets. Now, since each edge, $e^{\prime}$ or $e^{\prime\prime}$ is associated with a unique agent, an invocation of ”Protocol MYN” (Phase 1, Sec. 4.1) is required to inform the correct agent associated with a particular edge so that the agent can schedule the particular edge in the next step if required.
@@
    Checking for Termination: To determine whether the algorithm is ready for termination, each agent first evaluates the $change$ value for the current phase. Each agent then communicates with all other agents to collect their $change$ values and compute the binary product of these $n$ values. This process requires the application of Protocol MYN repeatedly (at least $D$ times) to ensure that every $change$ value is propagated to all $n$ agents. Given the importance of this step, which may be performed multiple times during iterations, we first establish a communication framework to propagate the binary product across all $n$ agents effectively.
    To achieve this, the agents construct a BFS tree of the graph, rooted at the node with the minimum ID agent. This BFS tree construction step is a self-independent one-time process and can be completed either just before the first iteration or at the beginning of the decomposition algorithm. During the iterations, if the binary product of the $change$ values results in $1$ (indicating that all agents have the same binary product value), the agents terminate the algorithm. If not, the agents repeat Phase 3 until the product equals $1$. Details of the BFS tree construction mechanism are provided in the subsequent section.
Creating a BFS Tree for Communication
As described earlier, to ensure the termination of the Truss Decomposition algorithm, the $n$ agents must communicate to verify that no further edges are scheduled for upgrading and that no additional iterations are required. This consensus information must be disseminated across all agents to confirm the completion of the algorithm. However, due to the constraints of local communication, propagating such information globally becomes the most time-intensive part of the algorithm.
To achieve termination through a global consensus, the agents must repeatedly execute Protocol MYN for at least $D$ rounds, where $D$ is the graph’s diameter, each time an edge is flagged for further upgrading. The algorithm terminates only after confirming that all agents reach the state $change=1$, indicating that no further computation is needed. This process intuitively adds a multiplicative factor of at least $O(D\Delta\log\lambda)$ rounds to the time complexity for each iteration (at-most $m$ times in the worst case), where $m$ is the number of edges.
The complexity of this process can be significantly reduced if, instead of applying Protocol MYN repeatedly, the agents construct a BFS tree rooted at the minimum ID agent, allowing for more efficient communication across the network.
In this subsection, we present an algorithm that constructs a BFS tree for the graph $G$, rooted at the agent with the minimum ID, denoted as $r^{\star}$. The primary purpose of this BFS tree is to compute a global value $x=f(x_{1},x_{2},x_{3},\dots,x_{n})$ from the individual values $x_{i}$ provided by each agent $r_{i}$, and then disseminate the computed value $x$ to all agents in the network in minimum amounts of time. To facilitate the construction of the BFS tree, each agent is equipped with the following variables.
  • $\bullet$
    $r_{i}.parent$, $r_{i}.child-$ used for storing respective $child$ and $parent$ pointers for $r_{i}$. We recall that each node $u$ with degree $\delta_{u}$ has a local port-numbering =$=\{0,1,\dots,\delta_{u}-1\}$. The $r_{i}.parent$ pointer stores the port number that connects $u$ to its parent agent. Similarly, the $child$ pointers store the port numbers that connect the node $u$ to its children. An agent has at-most $\Delta$ child pointers. Initially, $r_{i}.parent,r_{i}.child=\perp$. The child pointers uses $O(\Delta\log\Delta)$ bits overall, which is absorbed in the overall memory requirement of $O(\Delta\log n)$ bits for the truss decomposition.
  • $\bullet$
    $r_{i}.level-$ denotes the level assigned to an agent relative to the tree that $r_{i}$ is part of. Initially, $r_{i}.level=-1$
  • $\bullet$
    $r_{i}.visitor$ a variable set to $1$, if $r_{i}$ is a visitor to another node from its own node. When $r_{i}$ is at its home node, the $visitor$ value is 00.
The algorithm starts by electing a leader agent, which then coordinates the construction of a BFS tree of the graph $G$. After the leader is chosen, the agents proceed to build the BFS tree incrementally, level by level. To ensure proper synchronization between levels, the agents utilize the knowledge $\Delta$, the maximum degree of the graph. The algorithm operates over $O(D\Delta)$ rounds, with each level of the BFS tree assigned $2\Delta$ rounds for completion.
Step 1: Electing a Leader
To elect a leader, each agent $r_{i}$ first sends its ID to its neighbours. Each agent now computes the minimum of the IDs received from neighbouring agents and sends this minimum ID to its neighbours. Each round of communication with neighbours is performed using Procedure MYN. This process continues for $D$ rounds, with each communication requiring $O(\Delta\log\lambda)$ rounds. So, to elect a leader an $r_{i}$ does the following steps for $D$ rounds.
  1. 1.
    $r_{i}$ calculates the minimum IDs it has received from the neighbours.
  2. 2.
    $r_{i}$ sends this minimum ID to its neighbours with one execution of Procedure MYN.
Using Lemma 4.4, we can argue that after $d$ steps of communication, every agent $r_{i}$, now has the minimum ID of all the agents in its $d-hop$ neighbourhood. Since the diameter of the graph is $D$, we have the following lemma.
Lemma 5.4.
By the end of $O(D\cdot\Delta\log\lambda)$ rounds every agent has the knowledge of the least ID agent in the network.
The agents are now aware of the leader agent. We denote this agent as $r^{\star}$. $r^{\star}$ now initiates the construction of the BFS tree rooted at $r^{\star}$.
Step 2: Construction of BFS Tree
The leader agent $r^{\star}$ begins constructing a BFS tree, with itself as the root. The process proceeds level by level. First, $r^{\star}$ marks itself as level 00, initiating the algorithm by visiting all its neighbouring agents one by one. As $r^{\star}$ visits each neighbour, it marks them as its children, and these newly marked children also mark $r^{\star}$ as their parent simultaneously.
To explore the graph in levels, agents use the variable $level$. When $r^{\star}$ identifies its children (and they, in turn, recognize $r^{\star}$ as their parent), these agents update their level to $r^{\star}.level+1$. The children of $r^{\star}$ then begin exploring the next level in search of their own children. However, before they proceed, the children of $r^{\star}$ at level $1$ must ensure that all their siblings have been explored by $r^{\star}$. Only after this can agents at the next level begin exploring in parallel. This synchronization is crucial to correctly constructing the BFS tree. If any child prematurely starts extending the tree before all siblings have been explored, it could result in a tree with longer-than-necessary paths, violating the BFS structure. For example, in a complete graph, if the first discovered child immediately continues exploring its own children, the resulting tree could have $n-1$ levels, whereas the correct BFS tree should only have one level. To prevent this, the agents synchronize their actions by waiting for $2\Delta$ rounds.
Below is an outline of the algorithm for an agent $r_{i}$ during rounds $2k\Delta$ to $2(k+1)\Delta$:
  1. 1.
    Case 1: $r_{i}.level=k$
    • $\bullet$
      $r_{i}$ sets $r_{i}.visitor\leftarrow 1$ moves to its neighbours one by one using increasing port numbers (recall that each node has a local port numbering for every outgoing edge). Let one of its neighbours be $v$.
    • $\bullet$
      Sub-case 1: If $r_{i}$ meets a single new agent $r_{t}$ at $v$ with $level=-1$ (implying the agent has not been visited before), it marks $r_{t}$ as its child, and simultaneously $r_{t}$ sets $r_{i}$’s incoming port as its parent. Additionally, $r_{t}$ sets $r_{t}.level\leftarrow r_{i}.level+1$.
    • $\bullet$
      Sub-case 2: $r_{i}$ finds multiple agents at $v$
      • 1: Original resident present - If the original resident $r_{t}$ of $v$ (with $r_{t}.visitor=0$) is at $v$, $r_{t}$ accepts the request from the visitor agent with the lowest ID. If $r_{i}$ is that agent, it updates its child pointers accordingly. The newly discovered agent sets its parent accordingly and increments its level to $r_{i}.level+1$. If $r_{i}$ is not selected, it returns to its original node and proceeds to the next neighbour.
      • 2: Original resident absent - If the original resident is absent, all agents are visitors. In this case, the agent $r_{i}$ returns to its original home node.
    • $\bullet$
      Sub-case 3: If $r_{i}$ finds $v$ to be empty, $r_{i}$ returns to explore other neighbourhoods. An empty node indicates that the resident agent of that particular node is at the same level as $r_{i}$.
    • $\bullet$
      Sub-case 4: If $r_{i}$ encounters an agent $r_{t}$ that is already assigned a non-negative level, $r_{i}$ ignores $r_{t}$ and returns to its home node for further exploration.
    • $\bullet$
      If all the outgoing ports (neighbours) have been visited, $r_{i}$ sits at its node till the remaining time in $2(k+1)\Delta$ rounds has elapsed.
  2. 2.
    Case 2: $r_{i}.level\neq k$: In this case, $r_{i}$ does not take any action and remains idle at its current node during the $2\Delta$ rounds. However, if it is visited by a neighbouring agent (which could be a potential parent), $r_{i}$ behaves as described for $r_{t}$ in the previous scenario, responding appropriately to establish the parent-child relationship.
To complete the BFS tree construction, this process runs in phases of $2\Delta$, a $D$ number of times. Therefore, after the completion of $2\Delta D$ rounds, the algorithm terminates and the BFS construction from $r^{\star}$ is complete. Since the number of levels in a BFS tree cannot exceed the diameter $D$ of the graph, we have the following lemma.
Lemma 5.5 (BFS Tree Construction).
The $n$ agents construct a BFS tree for $G$ rooted at $r^{\star}$ in $O(D\Delta\log\lambda)$ rounds.
We now provide a mechanism that allows the root $r^{\star}$ to collect $n$ different values $x_{i}$ from each agent $r_{i}$ and compute a value $x=f(x_{1},x_{2},\cdots,x_{n})$ and then disseminate this value to all the agents in $G$.
  1. 1.
    Each agent $r_{i}$ sends its own value $x_{i}$ to its parent.
  2. 2.
    Each agent waits till it gets the array of values from all of its children, each appends its own value to the array and continues to forward this updated set of data to its parent.
  3. 3.
    Finally $r^{\star}$ after receiving all these $x_{i}$ values from its children and using its own value $x^{\star}$ computes the function $x=f(x_{1},x_{2},\cdots,x_{n})$
  4. 4.
    To send this value across the BFS tree, the agents use their $level$ number.
  5. 5.
    In the first $\Delta$ rounds, the agents with $level=1$ moves to its parent ($r^{\star}$) to collect the value of $x$ from $r^{\star}$.
  6. 6.
    In round $t$, each agent $r_{i}$ with $level=t$ similarly moves up to its parent to collect the value of $x$.
  7. 7.
    Finally, within $D$ rounds, this value of $x$ reaches every agent in the graph.
We therefore have the following lemma.
Lemma 5.6.
Each agent can learn the value of the function $f(x_{1},x_{2},\dots,x_{n})$, which is computed from $n$ inputs provided by the $n$ agents, in $O(D)$ rounds by utilizing the constructed BFS tree as the framework.
Therefore, using the methodology above, the agents can compute the binary product of the $change$ values, which is the function $f(x_{1},x_{2},\dots,x_{n})$ here, with $x_{i}$ representing the $change$ values of an agent $r_{i}$. In the following lemma, we now analyse the final time complexity of solving the Truss Decomposition Problem using the mobile agents.
Lemma 5.7.
Algorithm 5.1 takes $O(m\Delta\log\lambda+mD)$ rounds to terminate.
@@
Phase 1 and Phase 2 take $O(\Delta\log\lambda)$ rounds, as shown in Lemma 4.2 and Lemma 4.3. To elect a leader and construct the BFS tree, the agents need additional $O(D\Delta\log\lambda+D\Delta)$ rounds (Lemmas 5.45.5). In each execution of Phase 3, the agent needs to communicate with its neighbour to update its $trussness(h)$ values taking $O(\Delta\log\lambda)$ rounds. Finally, to check for termination, each agent propagates the $change$ value to every other agent, needing $O(D)$ rounds (Lemma 5.6).
At worst, the algorithm may update the $h-index$ value of one edge during each execution of Phase 3. Also, the up-gradation of the $h-index$ value of an edge is equivalent to the fact that the triangle count ($support)$ of that particular edge (and possibly some other edges) in $G$ has been decreased. Therefore, each time an $h-index$ of an edge is updated (it can only decrease), the graph $G$ loses at least one edge. Therefore, the algorithm may need to execute Phase 3, $m$ times before termination. Therefore Algorithm 5.1 takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda+D\Delta)+m\cdot(O(\Delta\log\lambda)+O(D))=O(m\Delta\log\lambda+mD)$ rounds to execute. ∎
@@
Let $G$ be an $n$-node arbitrary, simple, connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest agent ID $\lambda\in[0,n^{c}]$, where $c$ is constant, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, the Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta\log\lambda+mD)$ rounds with $O(\Delta\log n)$ bits of memory per agent.
@@ 5.2 Triangle Centrality
where, where $N(v)$ is the neighbourhood set of $v$, $N_{\mathbf{T}}(v)$ is the set of neighbours that are in triangles with $v$, and $N^{+}_{\mathbf{T}}(v)$ is the closed set that includes $v$. $\mathbf{T}(v)$ and $\mathbf{T}(G)$ denote the respective triangle counts based on $v$ and total triangle count in $G$. Mathematically, $N(v)=\{u:(u,v)\in E(G)\}$, $N_{\mathbf{T}}(v)=\{u\in N(v):N(u)\cap N(v)\neq\phi\}$ and $N^{+}_{\mathbf{T}}(v)=\{v\}\cup N_{\mathbf{T}}(v)$. Now, we outline our algorithm that computes $TC(v)$ for a node $v\in G$ using the mobile agents that run in these $3$ phases:
@@
    Computing $\mathbf{T}(v)$ and $\mathbf{T}(G)$ We use the triangle counting algorithm for mobile agents described in Section 4 to evaluate the number of triangles involving the node $v$, $\mathbf{T}(v)$ and the total number of triangles in $G$, $\mathbf{T}(G)$. The execution of this phase takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda)=O(D\Delta\log\lambda)$ rounds [Lemma 4.3,4.6]. With the completion of this phase, the agents begin the next phase. Note that if the count of the total number of triangles is available a priori, only $T(v)$ needs to be evaluated.
@@
    Computing $N(v)$, $N^{+}_{\mathbf{T}}(v)$ and $N_{\mathbf{T}}(v)$ $N(v)$, the neighbours of $v$ can be recorded in $O(\Delta\log\lambda)$ rounds by the agents (Lemma 4.2). After we allow the agents to record their neighbourhood in the first $O(\Delta\log\lambda)$ rounds, in the subsequent $O(\Delta\log\lambda)$ rounds, the agents can now meet their neighbours once again, with this time, learning about $N(u)$ for each neighbour agent $u$ of $v$. Now the agent from $v$ communicates with each $u\in N(v)$ to evaluate the set $N_{\mathbf{T}}(v)$. Similarly, the agents can also build the set $N^{+}_{\mathbf{T}}(v)$. Now, as the agent meets with each of its neighbours $u\in N(v)$, it cumulatively evaluates the sums $s=\sum_{u\in N^{+}_{\mathbf{T}}(v)}$ and $x=\sum_{u\in N(v)_{\mathbf{T}}(v)}$ simultaneously. Finally the sum $\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}=s-x+\mathbf{T}(v)$ is evaluated. This step takes $O(\Delta\log\lambda)$ rounds.